At a certain temperature, the solubility of strontium arsenate, Sr3(AsO4)2 , is 0.0660 g/L.

What is the
1
Expert's answer
2020-04-28T12:25:58-0400


The solubility of strontium arsenate is 0.0660 g/L

Its Molar Mass: 540.6984 g/mol

0.0660 / 540.6984 = 1,22x10-4 mol/L - molar solubility

[Sr2+] = 3 * 1,22x10-4 = 3,66x10-4 mol/L

[AsO43-] 2 * 1,22x10-4 = 2,44-4 mol/L



Ksp = (3,66x10-4)3 * (2,44x10-4)2 = 2,9x10-18